# On homology spheres of the trivial local equivalence class

### v1

@@ 2.3. Local equivalence in Seiberg–Witten Floer theory
  • the $\delta$-invariant [Frø10], the $\alpha$-, $\beta$-, and $\gamma$-invariants [Man16], and the $\underline{\delta}$- and $\overline{\delta}$-invariants [Sto17b] factor through $\mathfrak{CLE}$.
@@
Recall that $SWF(Y)$ recovers the $\operatorname{Pin}(2)$-equivariant Seiberg–Witten Floer homology of YY. [^2] Here, the Seiberg–Witten Floer homology means the $\operatorname{Pin}(2)$-equivariant Borel homology $\widetilde{H}_{*}^{\operatorname{Pin(2)}}(SWF(Y))$ of the space corresponding to $SWF(Y)$. By [LM18, Pan25], it turned out that this homology is isomorphic to the $\operatorname{Pin(2)}$-monopole Floer homology $\widecheck{HS_{*}}(Y)$ defined by [Lin18], based on [KM07]. Motivated from the facts that $\mathbb{H}^{-}(R)$ is isomorphic to $HF^{-}(Y)$ [Ném05], and hence to the Seiberg–Witten Floer homology [LM18, KLT20], Dai, Sasahira, and Stoffregen constructed the $\operatorname{Pin}(2)$-equivariant spectrum $\mathcal{H}(R)$ of the graded root $R$ of $Y$, and proved the following:
@@ 2.4. Filtered instanton Floer homology and $r_{s}$-invariants
We fix a Riemannian metric on $Y$ and an appropriate perturbation $\pi$ of $cs_{Y}$. Then we construct an infinite-dimensional analogue of Morse chain complex of $\widetilde{B}^{*}(Y)$ using the perturbed Chern–Simon functional $cs_{Y,\pi}$. The set $\widetilde{R}^{*}_{\pi}(Y)$ of the irreducible critical points of $cs_{Y,\pi}$ in $\widetilde{B}^{*}(Y)$, consists of flat connections, namely, those with curvature $F_{A}=0$. The integer-valued Floer index indind, which is an analogue of the ordinary Morse index, is also well-defined on $\widetilde{R}^{*}(Y)$. Then the instanton Floer chain complex is defined as:
@@
[…]
@@
3. Proof of Theorem 1.2
In this section, we prove Theorem 1.2. To do so, we first present explicit linear dependence relations of $h(Y_{1}(n))$, $h(Y_{2}(n))$ and $h(Y_{3}(n))$ in terms of $h(B(n))$ in the local equivalence group $\mathfrak{I}$ using Theorem 2.6. Then we prove linear independence of them in $\Theta_{\mathbb{Q}}^{3}$ by using Theorems 2.14 and 2.15.
We now describe the monotone graded subroots of $B(n)$ for $n\geq 1$ [DS19] (see also [Sto17a]) and the ones for the following infinite families [KŞ22] (see also [KŞ20]):
@@
  • $Y_{3}(n)=\Sigma(8n+1,12n+1,24n+5)$.
Theorem 3.1 (Theorem 1.1, [DS19]; Theorem 1.2, [KŞ22]).
@@
We next use Theorem 2.6 to decompose the monotone graded subroots above. Since $h:\Theta^{3}_{\mathbb{Z}}\to\mathfrak{I}$ is a homomorphism, for the case (1), we have:
@@
Next, we will show that all the Brieskorn spheres $B(n)$, $Y_{1}(n)$, and $Y_{2}(n)$ are linearly independent up to rational homology cobordism. In particular, the homology spheres in Theorem 1.2 are generated by the above.
Lemma 3.3.
The homology spheres below are linearly independent in $\Theta^{3}_{\mathbb{Q}}:$
$\big{\{}B(n)\big{\}}_{n\geq 0}\ \cup\ \big{\{}Y_{1}(n)\big{\}}_{n\geq 1}\ \cup\ \big{\{}Y_{2}(n)\big{\}}_{n\geq 1}.$
@@
Since every Brieskorn sphere bounds a negative definite $4$-manifold, by Theorem 2.11 and definition of the $r_{s}$-invariants,
$r_{0}(B(n))=r_{0}(Y_{1}(n))=r_{0}(Y_{2}(n))=\infty.$
By [KŞ20, Proposition 4.3], all of these homology spheres have negative definite plumbing graphs with central weights e=2e=-2. Thus, since R(Y)=2e3R(Y)=-2e-3 for any Seifert fibered sphere $Y$ [NZ85], we have:
@@
By Theorem 2.15, $r_{0}(-\Sigma(p,q,r))=\frac{1}{4pqr}$, provided $R(\Sigma(p,q,r))>0$. Thus, according to Theorem 2.14, it suffices to show that all the families of Brieskorn spheres have distinct product of the exponents. It is clear that $(2n+1)(4n+1)(4n+3)$, $(4n+1)(6n+2)(12n+1)$ and (4n1)(6n2)(12n1)(4n-1)(6n-2)(12n-1) are never the same as $2\cdot 3\cdot 5=30$, which implies that
$r_{0}(-B(0))$ $\neq r_{0}(-Y_{1}(m)),$
$r_{0}(-B(0))$ $\neq r_{0}(-Y_{2}(m)).$
For the family $B(n)$ with $n>0$, note that $(2n+1)(4n+1)(4n+3)$ is always odd. In contrast, we see that $(4m\pm 1)(6m\pm 2)(12m\pm 1)$ are always even. Thus, there are no pair of positive integers $(n,m)$ such that $(2n+1)(4n+1)(4n+3)=(4n\pm 1)(6m\pm 2)(12m\pm 1)$. Thus,
$r_{0}(-B(n))$ $\neq r_{0}(-Y_{1}(m)),$
$r_{0}(-B(n))$ $\neq r_{0}(-Y_{2}(m)).$
Finally, by dividing both of $(4m\pm 1)(6m\pm 2)(12m\pm 1)$ by $2$, define $f(a)=(4a+1)(3a+1)(12a+1)$ and g(b)=(4b1)(3b1)(12b1)g(b)=(4b-1)(3b-1)(12b-1). Then for any $a\geq 1$, we can see that $g(a)<f(a)<g(a+1)$. Since $g(b)$ is strictly increasing when $b\geq 1$, there cannot exist a pair of positive integers $(a,b)$ such that $f(a)=g(b)$, which implies that
$r_{0}(-Y_{1}(m))\neq r_{0}(-Y_{2}(k)).$
Now since $r_{0}(B(n))=r_{0}(Y_{1}(n))=r_{0}(Y_{2}(n))=\infty$, and $r_{0}(-B(n))$, $r_{0}(-Y_{1}(m))$ and $r_{0}(-Y_{2}(k))$ are finite and distinct, all of them are linearly independent in $\Theta_{\mathbb{Q}}^{3}$ by Theorem 2.14. ∎
For the same reason as in the proof of Lemma  3.3, we see that $R(Y_{3}(n))=1$. Since $(8n+1)(16n+1)(16n+3)\not\equiv(8n+1)(12n+1)(24n+5)$ mod $4$, we have that
$r_{0}(-B(4n))\neq r_{0}(-Y_{3}(n)).$
Thus, we have the following lemma as a consequence of Theorems 2.14 and 2.15:
Lemma 3.4.
The homology spheres below are linearly independent in $\Theta_{\mathbb{Q}}^{3}:$
$\big{\{}B(4n)\big{\}}_{n\geq 0}\ \cup\ \big{\{}Y_{3}(n)\big{\}}_{n\geq 1}.$
Now we are ready to prove Theorem 1.2.
Proof of Theorem 1.2.
Consider the homology spheres of the forms
$\big{\{}Y_{1}(n)\ \#\ -B(2n)\ \#\ B(n)\ \#\ -nB(0)\big{\}}_{n\geq 1}\cup\big{\{}Y_{2}(n)\ \#\ -B(2n-1)\ \#\ B(n)\ \#\ -nB(0)\big{\}}_{n\geq 1}\cdot$
By (1) and (2) in Lemma 3.2, we know that they lie in the kernel of the homomorphism $h:\Theta^{3}_{\mathbb{Z}}\to\mathfrak{I}$. Applying Lemma 3.3, we also see that they are linearly independent in $\Theta^{3}_{\mathbb{Q}}$, which completes the proof.
Consider the other family of homology spheres
$\{Y_{3}(n)\#-B(4n)\}_{n\geq 1}.$
[…]

### v2

@@ 2.3. Local equivalence in Seiberg–Witten Floer theory
  • the $\delta$-invariant [Fro10], the $\alpha$-, $\beta$-, and $\gamma$-invariants [Man16], and the $\underline{\delta}$- and $\overline{\delta}$-invariants [Sto17b] factor through $\mathfrak{CLE}$.
@@
Recall that $SWF(Y)$ recovers the $\operatorname{Pin}(2)$-equivariant Seiberg–Witten Floer homology of YY. [^2] Here, the Seiberg–Witten Floer homology means the $\operatorname{Pin}(2)$-equivariant Borel homology $\widetilde{H}_{*}^{\operatorname{Pin(2)}}(SWF(Y))$ of the space corresponding to $SWF(Y)$. By [LM18, Pan25], it turned out that this homology is isomorphic to the $\operatorname{Pin(2)}$-monopole Floer homology $\widecheck{HS_{*}}(Y)$ defined by [Lin18], based on [KM07]. Motivated from the facts that $\mathbb{H}^{-}(R)$ is isomorphic to $HF^{-}(Y)$ [Nem05], and hence to the Seiberg–Witten Floer homology [LM18, KLT20], Dai, Sasahira, and Stoffregen constructed the $\operatorname{Pin}(2)$-equivariant spectrum $\mathcal{H}(R)$ of the graded root $R$ of $Y$, and proved the following:
@@
In particular, if the associated homology sphere $Z_{Y}$ to a Seifert fibered sphere $Y$ is itself a Seifert fibered sphere, then the homology sphere $Y\#-Z_{Y}$ in $\ker(h)$ also lies in the kernel of $\Theta^{3}_{\mathbb{Z}}\rightarrow\mathfrak{LE}$.
@@ 2.4. Filtered instanton Floer homology and $r_{s}$-invariants
We fix a Riemannian metric on $Y$ and an appropriate perturbation $\pi$ of $cs_{Y}$. Then we construct an infinite-dimensional analogue of Morse chain complex of $\widetilde{B}^{*}(Y)$ using the perturbed Chern–Simons functional $cs_{Y,\pi}$. The set $\widetilde{R}^{*}_{\pi}(Y)$ of the irreducible critical points of $cs_{Y,\pi}$ in $\widetilde{B}^{*}(Y)$, consists of flat connections, namely, those with curvature $F_{A}=0$. The integer-valued Floer index indind, which is an analogue of the ordinary Morse index, is also well-defined on $\widetilde{R}^{*}(Y)$. Then the instanton Floer chain complex is defined as:
@@
Suppose $\theta_{Y}^{[s,r]}]\neq 0$. By definition of $i_{[s,r]}^{[s^{\prime},r^{\prime}]}$ and Lemma 2.12, $\theta_{Y}^{[s,r^{\prime}]}]$ vanishes for some $r^{\prime}<r$ and then $\theta_{Y}^{[s,r^{\prime\prime}]}]=0$ for any $r^{\prime\prime}\leq r^{\prime}$. This “birth-death” property motivates to define the $r_{s}$-invariant to measure when the obstruction class $\theta_{Y}^{[s,r]}]$ begins to vanish. The coefficient ring is taken as $\mathbb{Q}$, since one may want to divide $IW^{1}_{[s,r]}([\theta_{Y_{2}}^{[s,r]}])$ by the non-zero integer $|H_{1}(W;\mathbb{Z})|$ to obtain $\theta_{Y_{1}}^{[s,r]}]$ as in Theorem 2.11.
@@
3. Proof of Theorem 1.2 and Examples
In this section, we prove Theorem 1.2 which establishes conditions on a Seifert fibered homology sphere $Y$ to ensure that $Y$ and the associated homology sphere $Z_{Y}$ are different in $\Theta^{3}_{\mathbb{Q}}$, in terms of the product $\mathcal{E}(Y)$ of its exponents and the Fintushel–Stern $R$-invariant [FS85]. It is evident that the homology sphere $Y\#-Z_{Y}$ lies in $\operatorname{ker}(h)$, and hence provides a potential counterexample for Question 1.1. We restate Theorem 1.2 in a slightly stronger way which clearly implies the version in Section 1.
Theorem 1.2.
Let $\{Y(n)\}$ be a family of Seifert fibered sphere and let $Z(n)$ be the associated homology spheres to $Y(n)$ of the same image in $\mathfrak{I}$. If $R(Y(n))>0$, $\mathcal{E}(Y(n))>30$, and the integers $\mathcal{E}(Y(n))$ and $(2m+1)(4m+1)(4m+3)$ are all distinct, then $\{Y(n)\}\cup\{Z(n)\}$ are linearly independent in $\Theta_{\mathbb{Q}}^{3}$.
Proof.
Since $Z(n)$ is a linear combination of $\{B(m)\}$, it is enough to prove that $\{Y(n)\}$ and $\{B(m)\}$ are linearly independent in $\Theta_{\mathbb{Q}}^{3}$. Because every Seifert fibered sphere bounds a negative definite $4$-manifold, by Theorem 2.11 and the definition of the $r_{s}$-invariants,
$r_{0}(Y(n))=r_{0}(B(m))=\infty.$
By the assumption $R(Y(n))>0$, we can apply Theorem 2.15 to obtain
$r_{0}(-Y(n))=\frac{1}{4\mathcal{E}(Y(n))}\cdot$
Thus, under the assumptions that $\mathcal{E}(Y(n))$ and $(2m+1)(4m+1)(4m+3)=\mathcal{E}(B(m))$ are finite and distinct, by Theorem 2.14, the proof is complete. ∎
Now we present explicit examples of the associated homology spheres $Z_{1}(n)$, $Z_{2}(n)$, and $Z_{3}(n)$ to the following families $Y_{1}(n)$, $Y_{2}(n)$, and $Y_{3}(n)$ of certain Seifert fibered spheres:
@@
  • $Y_{3}(n)=\Sigma(8n+1,12n+1,24n+5)$,
respectively, in terms of $B(n)$ using Theorem 2.6. Then we prove linear independence of them in $\Theta_{\mathbb{Q}}^{3}$ by using Theorem 1.2.
We next describe the monotone graded subroots of $B(n)$ for $n\geq 1$ [DS19] (see also [Sto17]) and the ones for the families $Y_{1}(n)$, $Y_{2}(n)$, and $Y_{3}(n)$ [KS22] (see also [KS20]):
Theorem 3.1 (Theorem 1.1, [DS19]; Theorem 1.2, [KS22]).
@@
We next use Theorem 2.6 to decompose the monotone graded subroots above. Since $h:\Theta^{3}_{\mathbb{Z}}\rightarrow\mathfrak{I}$ is a homomorphism, for the case (1), we have:
@@
Therefore, the associated homology spheres $Z_{1}(n)$, $Z_{2}(n)$, and $Z_{3}(n)$ respectively to $Y_{1}(n)$, $Y_{2}(n)$, and $Y_{3}(n)$ are as follows:
  • $Z_{1}(n)=B(2n)\ \#\ -B(n)\ \#\ nB(0)$,
  • $Z_{2}(n)=B(2n-1)\ \#\ -B(n)\ \#\ nB(0)$,
  • $Z_{3}(n)=B(4n)$.
As an application of Theorem 1.2, we will show that $\{Y_{i}(n)\#-Z_{i}(n)\}_{n\geq 1}$ for $i=1,2,3$, which lie in the kernel of $h$, are linearly independent in $\Theta_{\mathbb{Q}}^{3}$.
Theorem 3.3.
The homology spheres
$\big\{Y_{1}(n)\#-Z_{1}(n)\big\}_{n\geq 1}\ \cup\ \big\{Y_{2}(n)\#-Z_{2}(n)\big\}_{n\geq 1}$
are linearly independent in $\operatorname{ker}(h)$ and $\Theta_{\mathbb{Q}}^{3}$.
@@
By [KS20, Proposition 4.3], all of these homology spheres have negative definite plumbing graphs with central weights e=2e=-2. Thus, since R(Y)=2e3R(Y)=-2e-3 for any Seifert fibered sphere $Y$ [NZ85], we have:
@@
To apply Theorem 1.2, it is enough to show that all $\mathcal{E}(B(n))$, $\mathcal{E}(Y_{1}(m))$ and $\mathcal{E}(Y_{2}(k))$ for any positive integers $n,m$ and $k$ are distinct. It is clear that $(2n+1)(4n+1)(4n+3)$, $(4m+1)(6m+2)(12m+1)$ and (4k1)(6k2)(12k1)(4k-1)(6k-2)(12k-1) are never the same as $2\cdot 3\cdot 5=30$.
Note that $(2n+1)(4n+1)(4n+3)$ is always odd. In contrast, we see that $(4m\pm 1)(6m\pm 2)(12m\pm 1)$ are always even. Thus, there are no pair of positive integers $(n,m)$ such that $(2n+1)(4n+1)(4n+3)=(4m\pm 1)(6m\pm 2)(12m\pm 1)$.
Finally, by dividing both of $(4m\pm 1)(6m\pm 2)(12m\pm 1)$ by $2$, define $f(a)=(4a+1)(3a+1)(12a+1)$ and g(b)=(4b1)(3b1)(12b1)g(b)=(4b-1)(3b-1)(12b-1). Then for any $a\geq 1$, we can see that $g(a)<f(a)<g(a+1)$. Since $g(b)$ is strictly increasing when $b\geq 1$, there cannot exist a pair of positive integers $(a,b)$ such that $f(a)=g(b)$, which implies that $\mathcal{E}(Y_{1}(m))\neq\mathcal{E}(Y_{2}(k))$, which completes the proof. ∎
On the other hand, consider $Y_{3}(n)$ and the associated one $Z_{3}(n)$. In this case, $Z_{3}(n)$ is also Seifert fibered, so by Theorem 2.8, we see that $\{Y_{3}(n)\#-Z_{3}(n)\}_{n\geq 1}$ have the trivial local equivalence class of $\operatorname{Pin(2)}$-equivariant Seiberg–Witten Floer stable homotopy type in $\mathfrak{LE}$. However, we also prove their linear independence in $\Theta_{\mathbb{Q}}^{3}$.
Theorem 3.4.
The homology spheres
$\big\{Y_{3}(n)\#-Z_{3}(n)\big\}_{n\geq 1}$
are linearly independent in $\operatorname{ker}(\Theta^{3}_{\mathbb{Z}}\rightarrow\mathfrak{LE})$ and $\Theta_{\mathbb{Q}}^{3}$.
Proof.
For the same reason as in the proof of Theorem  3.3, we find that $R(Y_{3}(n))=1$. Since
$(8n+1)(16n+1)(16n+3)\not\equiv(8m+1)(12m+1)(24m+5)\mod 4,$
we conclude that $\mathcal{E}(Y_{3}(n))$ and $\mathcal{E}(Z_{3}(m))$ are distinct and we can apply Theorem 1.2. ∎
Note that our examples in Theorem 3.3 give other families of homology spheres whose local equivalence classes (also of Seiberg–Witten Floer types for Theorem 3.4) are trivial, yet which are linearly independent in $\Theta^{3}_{\mathbb{Q}}$, as stated in Corollary 1.3.